OpenStudy (anonymous):

Intergral of 1/ sqrt(1+2x) dx ? Upper limit=12 Lower limit = 0.

OpenStudy (anonymous):

use u- substituion

OpenStudy (anonymous):

\[\int\limits_{0}^{12} 1\div \sqrt{1+2x} dx \] This is what it looks like .. im getting the answer 8 but its actually 4. please explain?

OpenStudy (anonymous):

u= 1 + 2x du= 2 dx

OpenStudy (anonymous):

1/2 intergral of 1/sqrt(u) du

OpenStudy (anonymous):

..still lnot understanding ..